# When Do Credal Sets Stabilize? Fixed-Point Theorems for Credal Set Updates

### v1

@@ 3. Main Results
In Caprio et al., 2024a , the authors propose Credal Bayesian Deep Learning (CBDL), a novel framework to carry out Bayesian deep learning. At training time, the scholar specifies prior and likelihood Finitely Generated Credal Sets (FGCSs), i.e. credal sets of prior and likelihood distributions having finitely many extreme elements. Then, the posterior credal set is derived by computing the Bayesian updating of all possible pairs of priors and likelihoods that are extrema of their respective credal sets. This can be formalized as follows.
@@
Let $\mathcal{P}\in\mathscr{C}\subset 2^{\Delta_{\Theta}}$ be our prior FGCS, where $\Theta$ is our parameter space, and denote the collection of its extreme elements as $\text{ex}\mathcal{P}$. Let $\mathcal{L}\subset 2^{\Delta_{\mathcal{Y}}}$ be our likelihood FGCS (where $\mathcal{Y}$ is the measurable sample space), and let $\text{ex}\mathcal{L}=\{L_{1},\ldots,L_{K}\}$, $K\in\mathbb{N}$.
Recall that a correspondence $\varphi:\Delta_{\Theta}\rightrightarrows\Delta_{\Theta}$ assigns to some $P\in\Delta_{\Theta}$ a subset $\varphi(P)\subset\Delta_{\Theta}$, and that $\varphi(A)\coloneqq\cup_{P\in A}\varphi(P)$, for all $A\subset\Delta_{\Theta}$ (Aliprantis and Border,, 2006, Section 17.1).
The updating procedure prescribed by CBDL for $\mathcal{P}$, then, can be written as a function $f:\mathscr{C}\rightarrow\mathscr{C}$ such that
$\mathcal{P}\mapsto f(\mathcal{P})\coloneqq\text{CH}\left(\varphi(\text{ex}\mathcal{P})\right),$ (1)
where $\text{CH}(\cdot)$ denotes the convex hull operator, and $\varphi:\Delta_{\Theta}\rightrightarrows\Delta_{\Theta}$ is a correspondence
$\text{ex}\mathcal{P}\ni P^{\text{ex}}\mapsto\varphi(P^{\text{ex}})\coloneqq\{g_{1}(P^{\text{ex}}),\ldots,g_{K}(P^{\text{ex}})\},$
where $K\in\mathbb{N}$, and for $k\in\{1,\dots,K\}$
$g_{k}(P^{\text{ex}})(A)\coloneqq\frac{P^{\text{ex}}(A)L_{k}(E\mid A)}{L_{k}(E)}$ (2)
@@
for measurable sets $A\subset\Theta$, and measurable $E\subset\mathcal{Y}$ such that $L_{k}(E)\neq 0$. Here, of course, the function $g_{k}$ computes the Bayesian updating of $P^{\text{ex}}$ via likelihood $L_{k}\in\text{ex}\mathcal{L}$. Let $\mu$ be a $\sigma$-finite dominating measure on $\Theta$. For $P^{\mathrm{ex}}$ with density $p^{\mathrm{ex}}$ and likelihood kernel $\ell_{k}(E\mid\theta)$, set $P^{\mathrm{ex}}(A)=\int_{A}p^{\mathrm{ex}}(\theta)\mu(\text{d}\theta)$, $L_{k}(E)=\int_{\Theta}\ell_{k}(E\mid\theta)p^{\mathrm{ex}}(\theta)\mu(\text{d}\theta)$, and define the prior-conditional evidence
@@
Fix the measurable set $E\subset\mathcal{Y}$ representing the gathered evidence. If the following three hold,
@@
    For all $P^{\text{ex}}\in\text{ex}\mathcal{P}$, $p^{\text{ex}}\in\mathscr{L}^{1}(\Theta,\mu)$. For all $L_{k}\in\text{ex}\mathcal{L}$, $\ell_{k}(E\mid\cdot)$ is continuous,
@@
then function $f$ is Hausdorff-continuous.
We can then apply Theorem 1 to show that the CBDL updating technique for a prior FGCS $\mathcal{P}$ defined in Caprio et al., 2024a (, Algorithm 1) has (at least) a fixed point.    $\triangle$
The next corollary shows that if $f$ satisfies something slightly stronger than continuity, namely being an “almost” contraction, then its fixed point is unique.
@@
Let $\mathcal{X}$ be a compact metric space, and endow $\Delta_{\mathcal{X}}$ with the weak topology. Consider a function $f:\mathscr{C}\rightarrow\mathscr{C}$ that represents our updating technique. If $f$ is an “almost” contraction, i.e.
@@
Proof.
From the proof of Theorem 1, we know that, under our assumptions, $(\mathscr{C},d_{H})$ is a Hausdorff compact metric space. The result, then, follows from Edelstein’s fixed point theorem (Aliprantis and Border,, 2006, Theorem 3.52). ∎
Although interesting, Corollary 2.1 is not enough to ensure that the sequence $(\mathcal{P}_{n})$ of successive updating of an initial credal set $\mathcal{P}_{0}\in\mathscr{C}$, where $\mathcal{P}_{n+1}=f(\mathcal{P}_{n})$, for all $n\in\mathbb{Z}_{+}$—called the orbit of $\mathcal{P}_{0}$ under ff—converges in the Hausdorff metric $d_{H}$ to a unique limiting credal set $\mathcal{P}_{\star}$. For this to hold, we need to require a somewhat stronger condition. [^4] And implicitly assume that $\mathcal{P}_{n}\neq\emptyset$, for all $n\in\mathbb{Z}_{+}$.
Corollary 2.2 (Convergence of $(\mathcal{P}_{n})$).
Let $\mathcal{X}$ be a compact metric space, and endow $\Delta_{\mathcal{X}}$ with the weak topology. Consider a function $f:\mathscr{C}\rightarrow\mathscr{C}$ that represents our updating technique. If $f$ is such that
@@
Proof.
From the proof of Theorem 1, we know that, under our assumptions, $(\mathscr{C},d_{H})$ is a compact metric space, which implies that it is also complete. Our claim, then, follows from the Boyd-Wong fixed point theorem (Boyd and Wong,, 1969, Theorem 1). ∎

### v2

@@ 3.1.1. Credal Bayesian Deep Learning (CBDL)
In Caprio et al., 2024a , the authors propose Credal Bayesian Deep Learning (CBDL), a novel framework for Bayesian deep learning. At training time, the practitioner specifies prior and likelihood Finitely Generated Credal Sets (FGCSs): The prior FGCS on the network parameters and the likelihood FGCS to possibly encode different network architectures. Then, the posterior credal set is derived by computing the Bayesian updating of all possible pairs of priors and likelihoods that are extrema of their respective credal sets. This can be formalized as follows.
@@
Let $\mathcal{P}\in\mathscr{C}\subset 2^{\Delta_{\Theta}}$ be our prior FGCS, where $\Theta$ is our parameter space; Denote the collection of its extreme elements as $\text{ex}\mathcal{P}$, and its members by $P^{\text{ex}}$. Let $\mathcal{L}\subset 2^{\Delta_{\mathcal{Y}}}$ be our likelihood FGCS (where $\mathcal{Y}$ is the measurable sample space), and let $\text{ex}\mathcal{L}=\{L_{1},\ldots,L_{K}\}$, $K\in\mathbb{N}$.
@@
Let $\mu$ be a $\sigma$-finite dominating measure (see Appendix A.3) for $\text{ex}\mathcal{P}$ on $\Theta$. For $P^{\mathrm{ex}}$ with density $p^{\mathrm{ex}}$ with respect to $\mu$ and likelihood kernel $\ell_{k}(E\mid\theta)$ (see Appendix A.4), set $P^{\mathrm{ex}}(A)=\int_{A}p^{\mathrm{ex}}(\theta)\mu(\text{d}\theta)$, $L_{k}(E)=\int_{\Theta}\ell_{k}(E\mid\theta)p^{\mathrm{ex}}(\theta)\mu(\text{d}\theta)$, and define the prior-conditional evidence
@@
Recall that a correspondence $\varphi:\Delta_{\Theta}\rightrightarrows\Delta_{\Theta}$ assigns to some $P\in\Delta_{\Theta}$ a subset $\varphi(P)\subset\Delta_{\Theta}$, and that $\varphi(A)\coloneqq\cup_{P\in A}\varphi(P)$, for all $A\subset\Delta_{\Theta}$ (Aliprantis and Border,, 2006, Section 17.1). The updating procedure prescribed by CBDL for $\mathcal{P}$, can be written as a function $f:\mathscr{C}\rightarrow\mathscr{C}$ such that
(1) $\mathcal{P}\mapsto f(\mathcal{P})\coloneqq\text{CH}\left(\varphi(\text{ex}\mathcal{P})\right),$
where $\varphi:\Delta_{\Theta}\rightrightarrows\Delta_{\Theta}$ is a correspondence $\text{ex}\mathcal{P}\ni P^{\text{ex}}\mapsto\varphi(P^{\text{ex}})\coloneqq\{g_{1}(P^{\text{ex}}),\ldots,g_{K}(P^{\text{ex}})\},$ where $K\in\mathbb{N}$, and for $k\in\{1,\dots,K\}$,
(2) $g_{k}(P^{\text{ex}})(A)\coloneqq\frac{P^{\text{ex}}(A)L_{k}(E\mid A)}{L_{k}(E)}$
for measurable sets $A\subset\Theta$, and measurable $E\subset\mathcal{Y}$ such that $L_{k}(E)\neq 0$. Here, of course, the function $g_{k}$ computes the Bayesian updating of $P^{\text{ex}}$ via likelihood $L_{k}\in\text{ex}\mathcal{L}$. The following is proven in Appendix B.2.
@@
Fix the measurable evidence set $E\subset\mathcal{Y}$. If the following three hold,
@@
    For all $P^{\text{ex}}\in\text{ex}\mathcal{P}$, $p^{\text{ex}}\in\mathscr{L}^{1}(\Theta,\mu)$, that is, $\int_{\Theta}|p^{\text{ex}}(\theta)|\mu(\text{d}\theta)<\infty.$ Futhermore, for all $L_{k}\in\text{ex}\mathcal{L}$, $\ell_{k}(E\mid\cdot)$ is continuous,
@@
then the function $f$ is Hausdorff-continuous.
We can then apply Theorem 1 to show that the CBDL updating technique for a prior FGCS $\mathcal{P}$ defined in Caprio et al., 2024a (, Algorithm 1) has (at least) a fixed point. A similar result holds also for imprecise Bayesian continual learning (Lu et al.,, 2025), a continual learning (CL) method in which successive updates of credal sets are used to address CL under specific trade-offs.
3.2. Uniqueness of a Fixed Point
The next corollary shows that if $f$ satisfies a strong form of non-expansiveness—namely it is strictly distance-decreasing in the Hausdorff metric—then its fixed point is unique. It is proved in Appendix B.3
@@ 3.2. Uniqueness of a Fixed Point
Let $\mathcal{X}$ be a compact metric space, and endow $\Delta_{\mathcal{X}}$ with the weak topology. Consider a function $f:\mathscr{C}\rightarrow\mathscr{C}$ that represents our updating technique. If $f$ satisfies
@@
3.3. Convergence of the Orbit to the Unique Fixed Point
Although interesting, Corollary 2.1 is not enough to ensure that the sequence $(\mathcal{P}_{n})$ of successive updating of an initial credal set $\mathcal{P}_{0}\in\mathscr{C}$, where $\mathcal{P}_{n+1}=f(\mathcal{P}_{n})$, for all $n\in\mathbb{Z}_{+}$—called the orbit of $\mathcal{P}_{0}$ under ff—converges in the Hausdorff metric $d_{H}$ to a unique limiting credal set $\mathcal{P}_{\star}$. For this to hold, we need to require a somewhat stronger condition, and assume that $\mathcal{P}_{n}\neq\emptyset$, for all $n\in\mathbb{Z}_{+}$. We next present a generalization to credal sets of Boyd-Wong’s fixed point theorem, whose classical version is provided in Appendix A.5. There, we also provide a remark justifying our choice to generalize Boyd–Wong to study the convergence of the orbit $(\mathcal{P}_{n})$: This approach allows us to state the result in maximal generality.
Theorem 3 (Convergence of Orbit $(\mathcal{P}_{n})$).
Let $\mathcal{X}$ be a Polish space, and endow $\Delta_{\mathcal{X}}$ with the weak topology. Consider a function $f:\mathscr{C}\rightarrow\mathscr{C}$ that represents our updating technique. If $f$ is such that